### v1

# Evaluating LLMs for Demographic-Targeted Social Bias Detection: A Comprehensive Benchmark Study

Abstract
Large-scale web-scraped text corpora used to train general-purpose AI models often contain harmful demographic-targeted social biases, creating a regulatory need for data auditing and developing scalable bias-detection methods. Although prior work has investigated biases in text datasets and related detection methods, these studies remain narrow in scope. They typically focus on a single content type (e.g., hate speech), cover limited demographic axes, overlook biases affecting multiple demographics simultaneously, and analyze limited techniques. Consequently, practitioners lack a holistic understanding of the strengths and limitations of recent large language models (LLMs) for automated bias detection. In this study, we present a comprehensive evaluation framework aimed at English texts to assess the ability of LLMs in detecting demographic-targeted social biases. To align with regulatory requirements, we frame bias detection as a multi-label task using a demographic-focused taxonomy. We then conduct a systematic evaluation with models across scales and techniques, including prompting, in-context learning, and fine-tuning. Using twelve datasets spanning diverse content types and demographics, our study demonstrates the promise of fine-tuned smaller models for scalable detection. However, our analyses also expose persistent gaps across demographic axes and multi-demographic targeted biases, underscoring the need for more effective and scalable auditing frameworks.
Keywords: Social bias, Bias detection, Prompting, Fine-tuning
[…]
Evaluating LLMs for Demographic-Targeted Social Bias Detection: A Comprehensive Benchmark Study
[…]
1 MPI-SWS and Saarland University, Saarbrücken, Germany
2 Huawei Technologies France S.A.S.U., Paris, France
3 Huawei Technologies Company Ltd., Shenzhen, China
ayanm@mpi-sws.org, {chenfeihao, jinghui.li, jasmine.xwang}@huawei.com
Abstract content
1.  Introduction
Large-scale web-scraped text corpora have significantly driven recent advances in general-purpose AI (GPAI) models. These corpora, however, are often minimally curated and can contain harmful social biases, e.g., hateful, toxic, or stereotypical content targeting various demographic axes Navigli et al. (2023). Consequently, training on such content can lead to unsafe model outputs and real-world harms Dodge et al. (2021); Vashney (2022).
Detecting demographic-targeted social biases in data has become both a regulatory and technical requirement Tabassi (2023); European Union (2024). For example, the EU’s GPAI Code of Practice highlights the importance of documenting demographic-centric harms in training corpora. Effective data-level mitigation strategies Gallegos et al. (2024) also depend on reliably identifying biased instances. While prior studies have examined biases in text corpora Kreutzer et al. (2022); Luccioni and Viviano (2021); Dodge et al. (2021), they typically rely either on small-scale human annotations or on simplistic automated techniques. Manual inspection, however, poses psychological risks Steiger et al. (2021), and despite advances in LLMs, few works offer a comprehensive analysis of the state of the art in social bias detection.
Compared to systematic benchmarks that examine LLMs for biased generation Parrish et al. (2022); Sun et al. (2024), detailed studies of their potential as tools for detecting social biases in text remain underexplored. Existing work is often limited across one or more dimensions: focusing on narrow demographics Wang et al. (2024), restricting attention to a single type of content such as hate Mathew et al. (2021), analyzing only specific contexts like certain subreddits Kumar et al. (2024), or analyzing only few methods such as zero-shot prompting Sun et al. (2024). Moreover, most studies overlook biases that target multiple demographic axes simultaneously, e.g., intersectionality. These gaps motivate a structured and comprehensive study of the current capabilities of recent LLMs in detecting demographic-targeted social biases.
To address these gaps, we propose an evaluation framework for systematically assessing recent LLMs in detecting demographic-targeted social biases in English text. Our findings reveal both the potential of fine-tuned smaller LLMs and existing performance disparities across demographics and multi-targeted cases. By establishing a structured foundation, this work provides practical guidance for regulatory compliance and safer AI deployment. Our main contributions are as follows.
  1. 1.
    We develop a demographic-focused taxonomy for social bias detection for better alignment with anti-discrimination principles and multi-axis targeted bias detection.
  2. 2.
    We frame bias detection as a multi-label task and build a detection testbed that supports prompting, in-context learning, and fine-tuning.
  3. 3.
    We adapt twelve widely used English datasets covering diverse content types and demographics targeted by social biases.
  4. 4.
    We conduct an extensive empirical evaluation assessing both detection accuracy and performance disparities across demographic axes.
[…]
2.  Related work
Bias in LLMs. Several works have evaluated biases in LLMs, independently analyzing content types like stereotypes Nadeem et al. (2021); Parrish et al. (2022) and hate/toxic content Gehman et al. (2020). Recently, Li et al. (2023) also studied the fairness of ChatGPT in binary decision-making. Several benchmarks also analyzed stereotype and toxic characteristics in generations of recently developed LLMs Wang et al. (2023); Sun et al. (2024); Wang et al. (2024).
Bias detection with LLMs. Prior work explored using LLMs in hate-speech moderation Barikeri et al. (2021); Mathew et al. (2021); Kumar et al. (2024); Zhan et al. (2025) or domain-specific bias detection Raza et al. (2024). Recent work Sun et al. (2024); Wang et al. (2024) also benchmarked prompt-based bias detection. However, they are restricted to zero-shot settings, cover limited demographics and data. In contrast, we systematically evaluate across multiple LLM-based methods, a broader range of demographics, and holistically analyze across different content types.
Bias analysis of corpora. Other work has directly analyzed large text corpora. Kreutzer et al. (2022) employed human surveys on a small web-crawled subset to assess multilingual quality and offensive content. Lexicon-based approaches have been used to detect opinion biases in Wikipedia Hube and Fetahu (2018). Luccioni and Viviano (2021) subsampled Common Crawl to study sexual and hateful content using n-grams, BERT, and logistic regression, while Dodge et al. (2021) analyzed C4, linking sentiment toward racial groups to biased QA outcomes. Although these studies provide valuable insights, they rely on relatively simplistic detection methods. In contrast, our work offers a systematic evaluation of state-of-the-art LLMs for social bias detection, providing deeper insights that can complement and extend prior analyses of large-scale corpora.
LLM guardrails. LLMs have also been explored as guardrails for GPAI systems Markov et al. (2023); Inan et al. (2023); Chen et al. (2024a); Zeng et al. (2024), primarily to mitigate harmful user prompts and model-generated outputs. While effective for moderating AI systems, these models are not designed for systematically identifying biases in raw text. As we later show, they fail to capture subtle social biases in texts, highlighting the need for dedicated evaluations and methods.
3.  Framework setup
Dataset Data Bias Taxonomy Coverage Content Samples
GEN SO DIS AGE RAC NAT REL SES PHY UNB
BBQ Parrish et al. (2022) Stereo 7843
BEC-Pro Bartl et al. (2020) Occup 2580
CrowS-pairs Nangia et al. (2020) Stereo 1278
HateXplain Mathew et al. (2021) Hate/Tox 5437
ImplicitHate ElSherief et al. (2021) Hate/Tox 2876
RedditBias Barikeri et al. (2021) Stereo 9087
SBIC Sap et al. (2020) Hate/Tox 5243
StereoSet Nadeem et al. (2021) Stereo 1417
ToxiGen Hartvigsen et al. (2022) Hate/Tox 2096
WinoBias Zhao et al. (2018) Occup 3168
Winogender Rudinger et al. (2018) Occup 240
Winoqueer Felkner et al. (2023) Stereo 5516
Table 1: Our adapted datasets, covering social-bias taxonomy and contents: stereotypes (Stereo), gender-occupation bias (Occup), and hate-toxicity (Hate/Tox). Empty cells: not covered; ✓: demographic target covered; ✥: involved in multi-axis targeted biases; gray cells: coverage after adaptations (Section 3.2).
Bias Type Labels Example
Unbiased UNB its always a mistaken identity because they think every black person is a criminal.Barikeri et al. (2021)
Single-axis RAC what are you talking about arabs are rich af. they got arab moneyBarikeri et al. (2021)
Multi-axis GEN,RAC I like my women like I like my black people DeadSap et al. (2020)
GEN,RAC interracism is a form of race war. it is black males waging sexual war of conquest and colonizing white wombs to make brown babiesElSherief et al. (2021)
Table 2: Representative examples showing biases targeting single and multiple demographic axes.
This section outlines the practical setup of our evaluation framework for analyzing the ability of LLMs to detect social biases in texts targeting different demographic groups. We first present the demographic-targeted taxonomy that underpins our framework, then describe how we adapt and integrate existing datasets for a holistic evaluation of bias detection. Finally, we detail the testbed we constructed to ensure comprehensive coverage of LLMs and detection methods.
3.1.  Demographic-targeted taxonomy
Prior evaluation studies are limited in scope: (i) they narrowly focus on specific content types (e.g., hate Mathew et al. (2021)), (ii) adopt narrow context-specific taxonomies (e.g., subreddits Kumar et al. (2024)), or (iii) consider only a small set of demographic axes Wang et al. (2024). Hence, we lack a holistic view of the current capabilities for detecting social biases Navigli et al. (2023), i.e., “prejudices, stereotypes, and discriminatory attitudes against certain groups” embedded in texts.
To address the limitations, our framework employs a demographic-centered taxonomy that identifies the demographic axes targeted by biased texts. This approach aligns directly with anti-discrimination regulations and governance measures European Commission (2025), while also enabling the study of multi-axis biases: cases where texts simultaneously target multiple groups, an aspect often overlooked. Concretely, our taxonomy spans nine axes with differing legal recognition:
  1. 1.
    Broad recognition: Gender identity (GEN), Sexual orientation (SO), Disability (DIS), Age (AGE), Race/ethnicity (RAC), Nationality (NAT), and Religion (REL), all widely protected Congress (1964); EU FRA (2018).
  2. 2.
    Narrow recognition: Socioeconomic status (SES) and Physical appearance (PHY), protected in certain regional frameworks Viprey (2002); Klose et al. (2025).
Any text not targeting one or more of these axes is considered “unbiased” (UNB) within our taxonomy. Each demographic axis serves as a prediction category, making the detection task twofold: (i) identify whether a text expresses demographic-targeted bias, and (ii) determine which demographics are targeted. Unlike prior benchmarks that treat bias detection as single-label or multi-class classification Mathew et al. (2021); Wang et al. (2024), our formulation supports multi-label prediction, capturing both single-axis (e.g., race only) and multi-axis (e.g., gender+race) biases (Table 2).
Our taxonomy also considers a broad range of content types encoding social biases through implicit harms Blodgett et al. (2020), including: i) Stereotype descriptions that stereotype, misrepresent, or disparage identities, ii) Occupation–gender associations that stereotype, erase, or exclude gender identities, and Hate or toxic content targeting demographics through toxicity, derogation, or dehumanization. However, by centering on demographic axes rather than surface content types, we enable systematic characterization of which demographics are harmed, allow for analysis of multi-axis cases, and align directly with anti-discrimination and AI governance frameworks European Union (2024).
3.2.  Adapting existing datasets
To enable comprehensive empirical evaluation in realistic settings, we adapt existing English datasets to align with our demographic-targeted social bias taxonomy. We surveyed widely used NLP datasets Gallegos et al. (2024), prioritizing diversity across demographic axes and harm types. Unlike prior benchmarks Wang et al. (2024), which often rely on fully GPT-generated categories (e.g., toxic text), we minimize synthetic data to reduce evaluation artifacts Koo et al. (2024); Maheshwari et al. (2024). [^1] The only exception is ToxiGen, although it has human-annotated GPT text, unlike Wang et al. (2024). Based on this review, we curated samples from twelve distinct datasets (Table 1).
Since some datasets were originally designed for evaluating generative biases, we applied minor adaptations similar to Wang et al. (2024). For StereoSet, we concatenate the context with the stereotype options. For BBQ, we construct text instances by pairing disambiguated contexts with answers. For CrowS-Pairs, we retain only the “more biased” sentences. For SBIC, we derive a single label using the majority vote across annotators. For ToxiGen, we label texts as biased only when human annotator scores indicate bias.
We also observed inconsistencies in how demographic groups are labeled across datasets, so we standardized labels for consistency. For instance, bias against “Arabs” or “Middle Eastern” identities, labeled inconsistently as RAC Nadeem et al. (2021) or REL Barikeri et al. (2021), is mapped to RAC, reserving REL for explicit religious references. Biases against national identities such as “Chinese” or “Mexican” are assigned to NAT. Bias against “Jewish” identity is annotated as both RAC and REL to reflect its ethnoreligious nature Litt (1961). We also improve regulatory alignment by disambiguating GEN and SO (e.g., transgender bias labeled as GEN) and denoting pregnancy-targeted bias under GEN rather than PHY Parrish et al. (2022). Labels outside our taxonomy (e.g., victim Sap et al. (2020)) are excluded.
The resulting dataset contains 46,781 entries, substantially larger than comparable benchmarks (e.g., 11,004 samples in Wang et al. (2024)). Biased instances are more prevalent (around 70%), with most targeting a single demographic axis and roughly 12% of biased instances targeting multiple axes simultaneously. Among demographic targets, GEN, RAC, SO, and REL are most common, while PHY is least prevalent. Multi-axis biases most frequently combine {GEN, SO} or {GEN, RAC}.
We split the dataset with 53% allocated to training and in-context setups and 47% to evaluation, reserving 10% of the training portion for hyperparameter tuning. To ensure robust evaluation, we remove test instances that are semantically very similar to training examples. Using all-MiniLM-L6-v2 embeddings with a cosine similarity threshold of 0.9, this deduplication removes 3,657 duplicates, producing a cleaner and more reliable benchmark.
$3.$  Methodological testbed
To ensure a comprehensive evaluation, we consider a testbed incorporating LLM-based detection methods that span both prompting and fine-tuning. Furthermore, we operationalize our testbed with a diverse suite of state-of-the-art, open-source, or open-weight LLMs spanning multiple paradigms and configurations.
3.3.1.  Prompting
Brown et al. (2020) demonstrated that instruction-tuned LLMs can effectively perform a variety of tasks through textual prompting in zero-shot scenarios. Following Inan et al. (2023); Palla et al. (2025), our evaluation framework employs policy-based prompting for bias detection. Specifically, the prompt includes a policy detailing the bias detection task and our demographic-based social bias taxonomy. To enhance the bias detection capabilities within this prompting framework, we also assess the benefits of incorporating few-shot examples for in-context learning. Specifically, we utilize the retrieval-augmented (RAG) framework Chen et al. (2024a), where an embedding model identifies the most relevant in-context examples for each input instance from the training/development set.
Models. We consider several instruction-tuned models ranging from 8B to 72B parameters, e.g., GLM-4 GLM et al. (2024), Llama-3.1 Dubey et al. (2024), and Qwen-2.5 Yang et al. (2024). We also analyze the guardrail model Llama Guard-3 Inan et al. (2023) to explore if such models could directly be applied for general text bias detection. To perform RAG-based in-context sample selection, we use the BGE-M3 Chen et al. (2024b) model.
3.3.2.  Fine-tuning
We also evaluate the effectiveness of aligning LLMs to the bias detection task through fine-tuning. The task is framed as multi-label prediction over the nine demographic axes. We solve it through sequence classification by attaching nine classifier nodes to a pre-trained LLM: to the [CLS] token for encoder-only models and to the final output token for decoder-only models.
Because detection must perform reliably across all demographic axes, our evaluation framework also explores the effectiveness of data reweighting Kamiran and Calders (2012) to address imbalances. Let $N$ denote the number of samples and $\mathcal{M}_{\phi}$ the model. For a given instance $i$, its labels $Y_{i}$ form a binary vector of length nine, where $Y_{i}^{m}=1$ if the demographic axis $m$ is targeted and 0 otherwise. The weighted loss is defined as:
$\begin{split}\mathcal{L}_{\text{FT}}=-&\frac{1}{9N}\sum_{i=1}^{N}\sum_{m=1}^{9}w_{i}\Big[\alpha_{m}Y_{i}^{m}\log\sigma_{m}(\mathcal{M}_{\phi}(d_{i}))\\ &+(1-Y_{i}^{m})\log\big(1-\sigma_{m}(\mathcal{M}_{\phi}(d_{i}))\big)\Big],\end{split}$
where $\alpha_{m}$ balances across demographic axes, and $w_{i}$ compensates for binary imbalances regarding biased and unbiased instances. All weights are derived from training data statistics.
Models. For encoder models, we consider RoBERTa Liu (2019) and DeBERTa He et al. (2020) and for decoder-only models we consider GPT-2 Radford et al. (2019). For each model, we consider various parameter scales where, across models, the parameters range from 125M to 1.5B.
3.4.  Evaluation metrics
Our comprehensive framework uses metrics capturing three dimensions: (i) distinguishing biased vs. unbiased text, (ii) accurate multi-label classification of bias types, and (iii) ensuring parity across demographic axes and multi-targeted texts.
Let $N$ be the number of evaluation instances. For each instance $i$, annotated labels are represented as $Y_{i}=\begin{pmatrix}Y_{i}^{m}\end{pmatrix}_{m=1}^{9}$ and model predictions as $\hat{Y}_{i}=\begin{pmatrix}\hat{Y}_{i}^{m}\end{pmatrix}_{m=1}^{9}$, where $Y_{i}^{m},\hat{Y}_{i}^{m}\in\{0,1\}$ denote whether axis $m$ is targeted (1) or not (0).
Binary bias detection. We reduce the multi-label task to a binary one by defining ground-truth labels $Y_{B_{i}}=1-\mathbb{I}[Y_{i}^{m}=0,,\forall m\in{1,\dots,9}]$, with predictions $\hat{Y}_{B_{i}}$ defined analogously. A value of $1$ indicates the presence of any bias, and 0 indicates none. On these binary labels, we report $F_{1}$, false positive rate (FPR), and false negative rate (FNR).
Multi-label bias detection. Alongside macro $F_{1}^{\text{M}}$ and micro $F_{1}^{\mu}$ scores, we report two multi-label measures Sorower (2010):
  • Exact Match Ratio: analyzing correctness of the full predicted label sets, $\operatorname{MR}=\tfrac{1}{N}\sum_{i=1}^{N}\mathbb{I}[\hat{Y}_{i}^{m}=Y_{i}^{m},\,\forall m]$, where higher scores are better.
  • Hamming Loss: analyzing the prediction’s partial coverage of label sets, $\operatorname{HL}=\tfrac{1}{9N}\sum_{i=1}^{N}\sum_{m=1}^{9}\mathbb{I}[Y_{i}^{m}\neq\hat{Y}_{i}^{m}]$, where lower scores are better.
Detection disparities. Our evaluation also examines whether LLMs not only detect social biases accurately but also exhibit systematic performance gaps across different demographic targets. Given $\mathcal{P}$ denotes FPR or FNR, we analyze disparities in the following scenarios:
  • Per-demographic. Following predictive fairness Hardt et al. (2016); Zafar et al. (2017), we compute the maximum absolute error gap, i.e., overall detection disparity across individual demographic axes: $\Delta_{\mathcal{P}}=\underset{m,m^{\prime}}{\text{max}}\big|\mathcal{P}_{m}-\mathcal{P}_{m^{\prime}}\big|$.
  • Multi-demographic. Inspired by Kearns et al. (2018), we measure if the models make systematically more errors in detecting biases that specifically target multiple axes simultaneously (e.g., {GEN,RAC}) relative to biases that target each constituent axis (e.g., only GEN or RAC). $\mathcal{G}^{\{m,m^{\prime}\}}_{\mathcal{P}}=\underset{x\in\{m,m^{\prime}\}}{\operatorname{max}}\big|\mathcal{P}_{\{m,m^{\prime}\}}-\mathcal{P}_{x}\big|$. This measure helps us understand if the FPR or FNR of multi-axis targeted biased instances is markedly higher, indicating potential blind spots for automated bias detection methods.
Method Model Setup Binary prediction Multi-label prediction Time
$F_{1}$ FPR FNR MR HL $F_{1}^{\mu}$ $F_{1}^{\text{M}}$
Prompting Llama Guard-3-8B 0-shot $68.94_{\pm 0.71}$ $0.184_{\pm 0.011}$ $0.440_{\pm 0.008}$ $372_{\pm 0.008} 0.0$$085_{\pm 0.001}$ $54.68_{\pm 0.80}$ $38.69_{\pm 1.62}$ 305
5-shot $75.16_{\pm 0.64}$ $0.192_{\pm 0.012}$ $0.358_{\pm 0.009}$ $485_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.66_{\pm 0.68}$ $46.24_{\pm 1.87}$ 354
10-shot $75.17_{\pm 0.64}$ $0.186_{\pm 0.011}$ $0.359_{\pm 0.008}$ $486_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.79_{\pm 0.69}$ $44.68_{\pm 1.82}$ 371
Llama-3.1-8B 0-shot $83.72_{\pm 0.45}$ $0.686_{\pm 0.013}$ $108_{\pm 0.005} 0.0$$046_{\pm 0.004}$ $0.202_{\pm 0.003}$ $49.17_{\pm 0.47}$ $36.01_{\pm 0.60}$ 307
5-shot $87.27_{\pm 0.40}$ $752_{\pm 0.012} 0.0$$023_{\pm 0.003}$ $0.411_{\pm 0.008}$ $0.140_{\pm 0.004}$ $62.19_{\pm 0.68}$ $44.58_{\pm 0.73}$ 359
10-shot $87.47_{\pm 0.40}$ $746_{\pm 0.013} 0.0$$021_{\pm 0.002}$ $0.501_{\pm 0.009}$ $0.127_{\pm 0.004}$ $64.69_{\pm 0.70}$ $45.96_{\pm 0.82}$ 378
GLM-4-9B 0-shot $83.65_{\pm 0.45}$ $769_{\pm 0.012} 0.0$$089_{\pm 0.005}$ $0.373_{\pm 0.008}$ $0.104_{\pm 0.002}$ $62.23_{\pm 0.60}$ $49.96_{\pm 1.60}$ 331
5-shot $87.10_{\pm 0.40}$ $774_{\pm 0.012} 0.0$$021_{\pm 0.003}$ $773_{\pm 0.007} 0.0$$036_{\pm 0.001}$ $85.95_{\pm 0.50}$ $73.43_{\pm 1.69}$ 351
10-shot $86.98_{\pm 0.41}$ $775_{\pm 0.012} 0.0$$023_{\pm 0.003}$ $782_{\pm 0.007} 0.0$$034_{\pm 0.001}$ $86.74_{\pm 0.48}$ $75.46_{\pm 1.68}$ 385
Llama-3.1-70B 0-shot $83.43_{\pm 0.49}$ $0.527_{\pm 0.014}$ $0.153_{\pm 0.006}$ $275_{\pm 0.008} 0.0$$098_{\pm 0.001}$ $66.46_{\pm 0.48}$ $55.66_{\pm 1.34}$ 545
5-shot $88.49_{\pm 0.38}$ $581_{\pm 0.014} 0.0$$046_{\pm 0.004}$ $657_{\pm 0.008} 0.0$$046_{\pm 0.001}$ $83.28_{\pm 0.42}$ $73.16_{\pm 1.36}$ 583
10-shot $88.82_{\pm 0.39}$ $557_{\pm 0.015} 0.0$$046_{\pm 0.004}$ $648_{\pm 0.008} 0.0$$047_{\pm 0.001}$ $83.08_{\pm 0.41}$ $75.07_{\pm 1.39}$ 591
Qwen-2.5-72B 0-shot $82.20_{\pm 0.47}$ $0.687_{\pm 0.013}$ $0.136_{\pm 0.006}$ $0.126_{\pm 0.006}$ $0.208_{\pm 0.004}$ $49.31_{\pm 0.50}$ $37.87_{\pm 0.55}$ 548
5-shot $87.24_{\pm 0.39}$ $551_{\pm 0.014} 0.0$$078_{\pm 0.005}$ $583_{\pm 0.008} 0.0$$065_{\pm 0.002}$ $77.33_{\pm 0.55}$ $60.44_{\pm 1.15}$ 584
10-shot $87.38_{\pm 0.41}$ $552_{\pm 0.014} 0.0$$075_{\pm 0.004}$ $600_{\pm 0.009} 0.0$$060_{\pm 0.002}$ $78.94_{\pm 0.52}$ $63.00_{\pm 1.19}$ 630
Fine-tuning RoBERTa-base unw. $90.80_{\pm 0.33}$ $299_{\pm 0.013} 0.0$$082_{\pm 0.004}$ $823_{\pm 0.006} 0.0$$026_{\pm 0.001}$ $89.15_{\pm 0.44}$ $81.30_{\pm 1.74}$ 13
rew. $92.04_{\pm 0.33}$ $328_{\pm 0.014} 0.0$$050_{\pm 0.004}$ $816_{\pm 0.007} 0.0$$027_{\pm 0.001}$ $89.33_{\pm 0.41}$ $82.14_{\pm 1.45}$ 13
RoBERTa-large unw. $91.20_{\pm 0.36}$ $221_{\pm 0.012} 0.0$$097_{\pm 0.005}$ $809_{\pm 0.007} 0.0$$027_{\pm 0.001}$ $88.43_{\pm 0.46}$ $82.75_{\pm 1.48}$ 36
[…]
DeBERTa-v2-XL unw. $92.70_{\pm 0.32}$ $203_{\pm 0.012} 0.0$$075_{\pm 0.004}$ $832_{\pm 0.006} 0.0$$024_{\pm 0.001}$ $89.86_{\pm 0.42}$ $82.94_{\pm 1.44}$ 104
rew. $93.84_{\pm 0.30}$ $225_{\pm 0.011} 0.0$$047_{\pm 0.004}$ $834_{\pm 0.006} 0.0$$024_{\pm 0.001}$ $90.35_{\pm 0.40}$ $83.31_{\pm 1.33}$ 102
DeBERTa-v3-large unw. $91.96_{\pm 0.34}$ $223_{\pm 0.012} 0.0$$083_{\pm 0.005}$ $825_{\pm 0.007} 0.0$$026_{\pm 0.001}$ $89.21_{\pm 0.44}$ $81.69_{\pm 1.66}$ 56
rew. $93.52_{\pm 0.30}$ $253_{\pm 0.012} 0.0$$044_{\pm 0.004}$ $814_{\pm 0.007} 0.0$$028_{\pm 0.001}$ $89.11_{\pm 0.42}$ $77.59_{\pm 1.29}$ 55
GPT-2-large unw. $89.36_{\pm 0.37}$ $0.295_{\pm 0.014}$ $0.110_{\pm 0.005}$ $795_{\pm 0.007} 0.0$$029_{\pm 0.001}$ $87.61_{\pm 0.46}$ $78.34_{\pm 1.58}$ 33
rew. $89.80_{\pm 0.35}$ $550_{\pm 0.014} 0.0$$029_{\pm 0.003}$ $815_{\pm 0.007} 0.0$$027_{\pm 0.001}$ $89.65_{\pm 0.40}$ $80.11_{\pm 1.49}$ 32
GPT-2-XL unw. $90.08_{\pm 0.37}$ $0.253_{\pm 0.013}$ $0.108_{\pm 0.005}$ $797_{\pm 0.007} 0.0$$029_{\pm 0.001}$ $87.81_{\pm 0.48}$ $79.67_{\pm 1.64}$ 82
rew. $91.20_{\pm 0.33}$ $426_{\pm 0.014} 0.0$$038_{\pm 0.003}$ $826_{\pm 0.006} 0.0$$025_{\pm 0.001}$ $90.11_{\pm 0.39}$ $82.67_{\pm 1.51}$ 82
Table 3: Bias detection using prompting (zero-shot or in-context) and fine-tuning (default unw. or reweighted rew. prediction loss). Binary indicates unbiased (negative) vs biased (positive) detection. Other measures are for multi-label bias prediction of bias targets. For MR and $F_{1}$ scores, higher is better; for HL, FPR, and FNR, lower is better. Time: median inference time in milliseconds.
Data Model Bin. $F_{1}$ MR HL
BBQ Llama-Guard-3-8B $16.79_{\pm 4.34}$ $0.082_{\pm 0.025}$ $0.103_{\pm 0.003}$
Llama-3.1-70B $73.70_{\pm 2.76}$ $962_{\pm 0.017} 0.0$$005_{\pm 0.002}$
DeBERTa-v2-XL $94.65_{\pm 1.39}$ $958_{\pm 0.018} 0.0$$006_{\pm 0.003}$
DeBERTa-v3-large $96.42_{\pm 1.14}$ $973_{\pm 0.014} 0.0$$004_{\pm 0.002}$
GPT2-XL $91.11_{\pm 1.76}$ $946_{\pm 0.021} 0.0$$007_{\pm 0.003}$
BEC-Pro Llama-Guard-3-8B $_{\pm 0.00} 0.000±0.0$$0_{\pm 0.000}$ $0.111_{\pm 0.000}$
Llama-3.1-70B $91.51_{\pm 2.09}$ $982_{\pm 0.015} 0.0$$002_{\pm 0.002}$
[…]
CrowS-Pairs Llama-Guard-3-8B $50.37_{\pm 4.25}$ $276_{\pm 0.037} 0.0$$086_{\pm 0.005}$
Llama-3.1-70B $95.19_{\pm 1.29}$ $640_{\pm 0.038} 0.0$$046_{\pm 0.006}$
DeBERTa-v2-XL $97.38_{\pm 0.96}$ $821_{\pm 0.030} 0.0$$026_{\pm 0.005}$
DeBERTa-v3-large $99.14_{\pm 0.53}$ $752_{\pm 0.034} 0.0$$039_{\pm 0.006}$
GPT2-XL $98.61_{\pm 0.66}$ $755_{\pm 0.033} 0.0$$040_{\pm 0.006}$
HateXplain Llama-Guard-3-8B $90.13_{\pm 0.92}$ $558_{\pm 0.021} 0.0$$067_{\pm 0.004}$
Llama-3.1-70B $91.51_{\pm 0.84}$ $418_{\pm 0.021} 0.0$$083_{\pm 0.004}$
DeBERTa-v2-XL $91.24_{\pm 0.88}$ $723_{\pm 0.018} 0.0$$039_{\pm 0.003}$
DeBERTa-v3-large $91.28_{\pm 0.89}$ $717_{\pm 0.020} 0.0$$042_{\pm 0.003}$
GPT2-XL $91.34_{\pm 0.86}$ $743_{\pm 0.019} 0.0$$036_{\pm 0.003}$
ImplicitHate Llama-Guard-3-8B $80.94_{\pm 1.85}$ $505_{\pm 0.026} 0.0$$066_{\pm 0.004}$
Llama-3.1-70B $99.03_{\pm 0.39}$ $657_{\pm 0.026} 0.0$$047_{\pm 0.004}$
DeBERTa-v2-XL $98.80_{\pm 0.42}$ $773_{\pm 0.022} 0.0$$037_{\pm 0.004}$
DeBERTa-v3-large $99.19_{\pm 0.35}$ $751_{\pm 0.023} 0.0$$039_{\pm 0.004}$
GPT2-XL $99.30_{\pm 0.32}$ $744_{\pm 0.022} 0.0$$039_{\pm 0.004}$
RedditBias Llama-Guard-3-8B $66.98_{\pm 1.58}$ $454_{\pm 0.020} 0.0$$070_{\pm 0.003}$
Llama-3.1-70B $79.85_{\pm 1.03}$ $716_{\pm 0.017} 0.0$$036_{\pm 0.002}$
DeBERTa-v2-XL $85.20_{\pm 1.06}$ $827_{\pm 0.014} 0.0$$023_{\pm 0.002}$
DeBERTa-v3-large $82.98_{\pm 1.07}$ $811_{\pm 0.015} 0.0$$025_{\pm 0.002}$
GPT2-XL $79.65_{\pm 1.11}$ $840_{\pm 0.014} 0.0$$020_{\pm 0.002}$
SBIC Llama-Guard-3-8B $80.02_{\pm 1.45}$ $431_{\pm 0.021} 0.0$$080_{\pm 0.003}$
Llama-3.1-70B $98.05_{\pm 0.41}$ $598_{\pm 0.020} 0.0$$056_{\pm 0.003}$
DeBERTa-v2-XL $99.65_{\pm 0.17}$ $754_{\pm 0.017} 0.0$$038_{\pm 0.003}$
DeBERTa-v3-large $99.38_{\pm 0.23}$ $698_{\pm 0.018} 0.0$$048_{\pm 0.003}$
GPT2-XL $99.49_{\pm 0.21}$ $725_{\pm 0.018} 0.0$$043_{\pm 0.003}$
StereoSet Llama-Guard-3-8B $35.93_{\pm 5.95}$ $190_{\pm 0.040} 0.0$$094_{\pm 0.005}$
Llama-3.1-70B $75.75_{\pm 3.44}$ $546_{\pm 0.050} 0.0$$056_{\pm 0.007}$
DeBERTa-v2-XL $77.16_{\pm 3.39}$ $770_{\pm 0.046} 0.0$$029_{\pm 0.006}$
DeBERTa-v3-large $79.26_{\pm 3.23}$ $744_{\pm 0.044} 0.0$$032_{\pm 0.006}$
GPT2-XL $74.47_{\pm 3.35}$ $749_{\pm 0.044} 0.0$$031_{\pm 0.006}$
ToxiGen Llama-Guard-3-8B $84.06_{\pm 2.73}$ $622_{\pm 0.045} 0.0$$054_{\pm 0.007}$
Llama-3.1-70B $82.23_{\pm 2.58}$ $659_{\pm 0.046} 0.0$$044_{\pm 0.007}$
DeBERTa-v2-XL $82.80_{\pm 2.67}$ $754_{\pm 0.041} 0.0$$037_{\pm 0.007}$
DeBERTa-v3-large $81.42_{\pm 2.79}$ $728_{\pm 0.041} 0.0$$038_{\pm 0.007}$
GPT2-XL $73.51_{\pm 3.02}$ $760_{\pm 0.042} 0.0$$033_{\pm 0.006}$
WinoBias-1 Llama-Guard-3-8B $82_{\pm 1.23} 0.0$$004_{\pm 0.007}$ $0.111_{\pm 0.001}$
Llama-3.1-70B $41.61_{\pm 5.12}$ $467_{\pm 0.063} 0.0$$060_{\pm 0.007}$
DeBERTa-v2-XL $91.77_{\pm 2.46}$ $951_{\pm 0.026} 0.0$$005_{\pm 0.003}$
DeBERTa-v3-large $92.98_{\pm 2.23}$ $972_{\pm 0.021} 0.0$$003_{\pm 0.002}$
GPT2-XL $60.34_{\pm 4.34}$ $852_{\pm 0.043} 0.0$$016_{\pm 0.005}$
WinoBias-2 Llama-Guard-3-8B $83_{\pm 1.32} 0.0$$004_{\pm 0.007}$ $0.111_{\pm 0.001}$
Llama-3.1-70B $49.59_{\pm 4.83}$ $581_{\pm 0.062} 0.0$$047_{\pm 0.007}$
DeBERTa-v2-XL $98.98_{\pm 0.89}$ $992_{\pm 0.011} 0.0$$001_{\pm 0.001}$
DeBERTa-v3-large $98.77_{\pm 0.98}$ $988_{\pm 0.013} 0.0$$001_{\pm 0.001}$
GPT2-XL $98.98_{\pm 0.84}$ $1.000_{\pm 0.000}$ $0.000_{\pm 0.000}$
WinoGender Llama-Guard-3-8B $_{\pm 0.00} 0.000±0.0$$0_{\pm 0.000}$ $0.111_{\pm 0.000}$
Llama-3.1-70B $63.33_{\pm 15.17}$ $623_{\pm 0.165} 0.0$$042_{\pm 0.018}$
DeBERTa-v2-XL $89.86_{\pm 7.98}$ $913_{\pm 0.100} 0.0$$010_{\pm 0.011}$
DeBERTa-v3-large $84.68_{\pm 9.29}$ $910_{\pm 0.097} 0.0$$010_{\pm 0.011}$
GPT2-XL $78.86_{\pm 10.71}$ $940_{\pm 0.076} 0.0$$007_{\pm 0.008}$
WinoQueer Llama-Guard-3-8B $92.09_{\pm 0.81}$ $829_{\pm 0.015} 0.0$$022_{\pm 0.002}$
Llama-3.1-70B $99.79_{\pm 0.14}$ $755_{\pm 0.017} 0.0$$028_{\pm 0.002}$
[…]
GPT2-XL $0_{\pm 0.00} 1.000±0.001$$_{\pm 0.001}$ $0.000_{\pm 0.000}$
Table 4: Detection performance (binary $F_{1}$, multi-label MR, HL) per constituent dataset for select models (prompt: 10-shot, fine-tune: rew. loss).
4.  Evaluating social bias detection
This section illustrates how our comprehensive evaluation study enables the practical assessment of LLM-based methods for detecting demographic-targeted social biases in text. Our analysis reveals both the strengths and current limitations of these approaches. For rigorous assessment of the overall results, we perform 1,000 bootstrap samples with replacement on the combined test set to compute 95% confidence intervals. Table 3 presents a detailed comparison of prompting and fine-tuning, reporting both binary performance (biased vs. unbiased) and multi-label categorization. We also report median inference time (in milliseconds) for each text instance. Moreover, for more fine-grained analysis, Table 4 reports bias detection performance of select prompted and fine-tuned LLMs for the twelve constituent datasets.
4.1.  Prompting methods
Our detailed results in Table 3 show how bias detection with prompting is highly sensitive to both in-context learning and model capacity.
In-context learning substantially improves detection. Across all models, providing few-shot examples consistently boosts binary $F_{1}$, lowers FNR, and increases multi-label metrics (MR, HL, $F_{1}^{\text{M}}$). Gains are significant with as few as five examples, while moving from five to ten shots yields only marginal improvements. However, inference time grows with the number of examples, highlighting the accuracy–efficiency tradeoff in prompting. Beyond the reported results, we also analyzed alternative in-context setups (in the appendix). We found that (i) RAG-based example selection outperforms random sampling, and (ii) alternative embeddings Youdao (2023) yield comparable results.
Model size and architecture strongly shape results. Larger models (e.g., Llama-70B, Qwen-72B) achieve higher binary and multi-label performance than smaller variants. Within model families, scale matters: Llama-70B outperforms Llama-8B across nearly all metrics. However, size alone is not decisive. GLM-4-9B rivals or surpasses larger Llama and Qwen models on multi-label metrics, and Llama-3.1-70B outperforms Qwen-2.5-72B despite similar scale. Larger models tend to reduce FPR but can increase FNR, reflecting greater sensitivity at the cost of more false negatives. Inference time rises steeply with model scale, from 350ms for 8B models to over 600ms for 70B+ models.
Guardrail models transfer poorly. Llama Guard, fine-tuned for moderation, consistently underperforms instruction-tuned models of similar size (e.g., Llama-3.1-8B), achieving the lowest binary $F_{1}$ and highest FNR. While effective at detecting overt hate speech, it fails to capture subtler biases like stereotypes.
Per-dataset analysis. From Table 4, we see the role of scale from the binary $F_{1}$ scores. The 70B Llama model outperforms smaller variants across most datasets. Llama-Guard shows lower overall binary $F_{1}$, but performs relatively well on hate/toxic content. It specifically achieves the highest score across all models on Toxigen, which is GPT-generated. This finding is consistent with its design as a GPAI guardrail. Multi-label metrics show that even larger models struggle to correctly identify bias types, especially for stereotyped contents, e.g., StereoSet and RedditBias.
Takeaway. Instruction-tuned LLMs with in-context learning and sufficient capacity provide the most effective prompting-based bias detection, although at the cost of increased inference time. We further show that guardrail-focused models alone are insufficient for holistic social bias identification.
4.2.  Fine-tuning methods
Our results in Table 3 show how the performance of fine-tuned LLM-based bias detectors is shaped by model size, architecture, and optimization strategy.
Fine-tuning substantially improves detection. Even small models, such as RoBERTa-base, surpass much larger prompting-only models (Llama-3.1-70B, Qwen-2.5-72B) on binary $F_{1}$ (above 90 vs. below 89) and multi-label metrics (MR, HL, micro $F_{1}^{\mu}$ and macro $F_{1}^{\text{M}}$). Fine-tuned models also achieve lower FNR and higher reliability in detecting biased content. Inference is far faster: RoBERTa completes batches in seconds, whereas prompting with 70B+ LLMs requires hundreds of seconds.
Architecture influences performance. Encoder models (RoBERTa, DeBERTa) consistently outperform decoder models (GPT-2), irrespective of scale. GPT-2-XL underperforms on binary and multi-label detection. In contrast, DeBERTa-v2-XL and RoBERTa-large achieve higher detection scores. Inference times also reflect architectural complexity: decoder models remain faster, whereas DeBERTa-v2-XL is particularly slow due to disentangled attention He et al. (2020).
Scaling improves detection. Within encoder families, larger variants (RoBERTa-large, DeBERTa-XL) achieve better detection results. Importantly, despite being the newer variant, DeBERTa-v3-large performs slightly worse than the larger but older DeBERTa-v2-XL. GPT-2 shows similar scaling trends within decoder models. Inference time increases with model size, reinforcing the tradeoff between accuracy and efficiency.
Loss reweighting has tradeoffs. Reweighted loss consistently improves binary FNR and macro $F_{1}^{\text{M}}$ (e.g., DeBERTa-v2-XL, RoBERTa-large, GPT-2-XL) by capturing subtle biases, but can raise FPR, particularly in decoder models. Effects are uneven: DeBERTa-v3-large shows reduced MR and macro $F_{1}^{\text{M}}$, suggesting reweighting may destabilize multi-label detection for some scenarios.
Per-dataset analysis. Table 4 shows how fine-tuned models achieve stronger binary detection across most datasets compared to prompting-based LLMs. Encoder models (DeBERTa) generally outperform decoder-only GPT-2, which remains competitive on many datasets but struggles with subtle stereotype cases, e.g, RedditBias and Winogender. For multi-label detection, DeBERTa-v2-XL shows consistently lower HL, indicating more accurate detection of demographic axes targeted.
Takeaway. Fine-tuned encoder models provide the most effective bias detection, outperforming prompting much larger models in both binary detection and identification of the correct bias types. Decoder-based models can perform well but show weaknesses on subtle stereotype datasets. Fine-tuning with reweighted loss improves recall, but may increase false positives, highlighting important tradeoffs that require consideration.
5.  Evaluating detection disparities
Method Model Setup Per-demographic disparity Multi-demographic targeted disparity
$\Delta_{\text{FNR}}$ $\Delta_{\text{FPR}}$ $\mathcal{G}^{\{\text{GEN,SO}\}}_{\text{FNR}}$ $\mathcal{G}^{\{\text{GEN,SO}\}}_{\text{FPR}}$ $\mathcal{G}^{\{\text{GEN,RAC}\}}_{\text{FNR}}$ $\mathcal{G}^{\{\text{GEN,RAC}\}}_{\text{FPR}}$
Prompting Llama Guard-3-8B 0-shot $510_{\pm 0.037} 0.0$$046_{\pm 0.005}$ $558_{\pm 0.019} 0.0$$020_{\pm 0.002}$ $537_{\pm 0.016} 0.0$$070_{\pm 0.004}$
5-shot $724_{\pm 0.031} 0.0$$045_{\pm 0.005}$ $776_{\pm 0.016} 0.0$$020_{\pm 0.002}$ $717_{\pm 0.014} 0.0$$074_{\pm 0.005}$
10-shot $756_{\pm 0.028} 0.0$$047_{\pm 0.005}$ $795_{\pm 0.015} 0.0$$019_{\pm 0.002}$ $718_{\pm 0.014} 0.0$$074_{\pm 0.004}$
Llama-3.1-8B 0-shot $0.605_{\pm 0.070}$ $0.424_{\pm 0.010}$ $0.548_{\pm 0.070}$ $0.278_{\pm 0.007}$ $428_{\pm 0.066} 0.0$$054_{\pm 0.010}$
5-shot $0.259_{\pm 0.085}$ $0.194_{\pm 0.009}$ $212_{\pm 0.074} 0.0$$096_{\pm 0.006}$ $112_{\pm 0.052} 0.0$$028_{\pm 0.006}$
10-shot $0.300_{\pm 0.104}$ $0.208_{\pm 0.009}$ $277_{\pm 0.077} 0.0$$089_{\pm 0.006}$ $144_{\pm 0.064} 0.0$$051_{\pm 0.008}$
GLM-4-9B 0-shot $0.603_{\pm 0.027}$ $0.428_{\pm 0.010}$ $0.582_{\pm 0.072}$ $0.187_{\pm 0.006}$ $0.264_{\pm 0.078}$ $0.281_{\pm 0.009}$
5-shot $378_{\pm 0.099} 0.0$$071_{\pm 0.005}$ $535_{\pm 0.074} 0.0$$095_{\pm 0.006}$ $0.334_{\pm 0.076}$ $0.101_{\pm 0.006}$
10-shot $349_{\pm 0.102} 0.0$$069_{\pm 0.005}$ $495_{\pm 0.075} 0.0$$097_{\pm 0.006}$ $0.318_{\pm 0.073}$ $0.103_{\pm 0.006}$
Llama-3.1-70B 0-shot $0.433_{\pm 0.074}$ $0.312_{\pm 0.009}$ $0.736_{\pm 0.064}$ $0.181_{\pm 0.006}$ $0.262_{\pm 0.079}$ $0.176_{\pm 0.007}$
5-shot $0.288_{\pm 0.105}$ $0.147_{\pm 0.007}$ $158_{\pm 0.071} 0.0$$75_{\pm 0.006} 0.0$$39_{\pm 0.042} 0.0$$070_{\pm 0.007}$
10-shot $0.274_{\pm 0.098}$ $0.176_{\pm 0.008}$ $164_{\pm 0.072} 0.0$$78_{\pm 0.007} 0.0$$88_{\pm 0.052} 0.0$$061_{\pm 0.006}$
Qwen-2.5-72B 0-shot $0.369_{\pm 0.020}$ $0.372_{\pm 0.010}$ $0.244_{\pm 0.076}$ $0.143_{\pm 0.007}$ $0.466_{\pm 0.076}$ $0.186_{\pm 0.010}$
5-shot $0.189_{\pm 0.024}$ $0.117_{\pm 0.008}$ $268_{\pm 0.075} 0.0$$052_{\pm 0.006}$ $109_{\pm 0.061} 0.0$$048_{\pm 0.007}$
10-shot $0.199_{\pm 0.050}$ $0.108_{\pm 0.006}$ $288_{\pm 0.076} 0.0$$63_{\pm 0.006} 0.0$$97_{\pm 0.062} 0.0$$037_{\pm 0.007}$
Fine-tuning RoBERTa-base unw. $490_{\pm 0.104} 0.0$$032_{\pm 0.004}$ $604_{\pm 0.071} 0.0$$029_{\pm 0.003}$ $549_{\pm 0.074} 0.0$$056_{\pm 0.004}$
rew. $185_{\pm 0.073} 0.0$$054_{\pm 0.005}$ $324_{\pm 0.072} 0.0$$058_{\pm 0.004}$ $251_{\pm 0.071} 0.0$$042_{\pm 0.005}$
RoBERTa-large unw. $307_{\pm 0.084} 0.0$$029_{\pm 0.004}$ $713_{\pm 0.067} 0.0$$027_{\pm 0.003}$ $548_{\pm 0.073} 0.0$$041_{\pm 0.004}$
rew. $192_{\pm 0.063} 0.0$$052_{\pm 0.005}$ $436_{\pm 0.077} 0.0$$036_{\pm 0.003}$ $373_{\pm 0.078} 0.0$$044_{\pm 0.004}$
DeBERTa-v2-XL unw. $312_{\pm 0.082} 0.0$$030_{\pm 0.004}$ $393_{\pm 0.077} 0.0$$027_{\pm 0.003}$ $564_{\pm 0.072} 0.0$$042_{\pm 0.004}$
rew. $208_{\pm 0.044} 0.0$$040_{\pm 0.004}$ $278_{\pm 0.072} 0.0$$034_{\pm 0.003}$ $305_{\pm 0.075} 0.0$$029_{\pm 0.004}$
DeBERTa-v3-large unw. $465_{\pm 0.107} 0.0$$026_{\pm 0.003}$ $625_{\pm 0.073} 0.0$$024_{\pm 0.003}$ $628_{\pm 0.061} 0.0$$033_{\pm 0.003}$
rew. $258_{\pm 0.089} 0.0$$053_{\pm 0.004}$ $388_{\pm 0.079} 0.0$$058_{\pm 0.004}$ $289_{\pm 0.074} 0.0$$038_{\pm 0.004}$
GPT-2-large unw. $483_{\pm 0.073} 0.0$$031_{\pm 0.004}$ $470_{\pm 0.070} 0.0$$043_{\pm 0.003}$ $779_{\pm 0.041} 0.0$$051_{\pm 0.004}$
rew. $271_{\pm 0.070} 0.0$$078_{\pm 0.006}$ $477_{\pm 0.078} 0.0$$084_{\pm 0.005}$ $261_{\pm 0.073} 0.0$$072_{\pm 0.005}$
GPT-2-XL unw. $367_{\pm 0.059} 0.0$$038_{\pm 0.004}$ $462_{\pm 0.075} 0.0$$031_{\pm 0.003}$ $602_{\pm 0.070} 0.0$$057_{\pm 0.004}$
rew. $300_{\pm 0.084} 0.0$$060_{\pm 0.005}$ $388_{\pm 0.071} 0.0$$051_{\pm 0.004}$ $299_{\pm 0.074} 0.0$$062_{\pm 0.005}$
Table 5: Detection disparity in terms of FPR and FNR (considering singular targets) and disparity for multi-label biased instances (targeting {GEN,SO}, {GEN,RAC})
We use our evaluation framework to examine potential disparities in social bias detection across models and setups with respect to targeted demographic axes. While the previous analysis provided a global view of model performance, this section focuses on systematic differences in how effectively models detect biases. We first analyze disparities for individual demographic axes. Next, owing to our multi-label setup, we evaluate model performances on instances targeting multiple axes simultaneously, highlighting current capabilities in detecting multi-targeted biases. We provide the comprehensive disparity analysis in Table 5.
5.1.  Per-demographic axis disparity
We assess systemic performance disparities using $\Delta_{\text{FNR}}$ and $\Delta_{\text{FPR}}$, which measure the maximum performance gaps across the nine social bias demographic target axes in our taxonomy.
Prompting shows large disparities. In zero-shot settings, models exhibit significant disparities. For instance, Llama-3.1-8B and GLM-4-9B exhibit $\Delta_{\text{FNR}}\approx 0.6$, $\Delta_{\text{FPR}}\approx 0.42$. Few-shot prompting reduces disparities (e.g., for Llama-3.1-8B, $\Delta_{\text{FNR}}$ drops to $\approx 0.26$), but performance remains uneven compared to fine-tuned models. Scaling improves parity: Llama-3.1-70B shows lower disparities than its 8B counterpart, and Qwen-2.5-72B achieves the strongest parity among prompting models, especially with few-shot examples.
Fine-tuning yields markedly lower disparities. Encoder models such as RoBERTa-large and DeBERTa-v2-XL reach $\Delta_{\text{FNR}}\approx 0.2$ and $\Delta_{\text{FPR}}\approx 0.03$, particularly with reweighted loss. Reweighting reduces FNR gaps but can slightly increase FPR gaps, indicating a tradeoff. Model architecture also matters: encoder models achieve far lower disparities than decoder-only GPT-2, and scaling further improves parity (e.g., RoBERTa-large outperforms RoBERTa-base).
Takeaway. Prompting, even with larger models and few-shot examples, shows substantial per-axis disparities. Fine-tuned models, particularly with reweighted loss, achieve more balanced performance, although notable gaps remain. In additional analyses (in the appendix), we examined $F_{1}$ scores across the nine demographic axes. We found that certain targets, such as NAT and PHY, consistently have lower detection accuracy, which contributes to the observed disparities.
5.2.  Multi-demographic disparity
We now analyze texts targeting multiple axes simultaneously, focusing on {GEN,SO} and {GEN,RAC}. We compare performance on these multi-axis instances to that on the instances that target only constituent single axes (e.g., only GEN or SO for {GEN,SO}).
Prompting improves with scale and few-shot examples. Few-shot prompting substantially narrows multi-axis gaps. For Llama-3.1-70B, $\mathcal{G}^{\{\text{GEN,SO}\}}_{\text{FNR}}$ drops from 0.736 (zero-shot) to 0.164 (10-shot), and $\mathcal{G}^{\{\text{GEN,RAC}\}}_{\text{FNR}}$ from 0.262 to 0.088. Larger models benefit more from in-context learning: Llama-3.1-70B outperforms Llama-3.1-8B, and disparities are generally higher for {GEN,SO} than {GEN,RAC}.
Fine-tuned models show persistent gaps. Despite strong per-axis parity, fine-tuned models underperform on multi-axis instances, reflecting gerrymandering Kearns et al. (2018) in their detection performance. For example, RoBERTa-large with reweighting achieves $\mathcal{G}^{\{\text{GEN,SO}\}}_{\text{FNR}}\approx 0.436$ and $\mathcal{G}^{\{\text{GEN,RAC}\}}_{\text{FNR}}\approx 0.373$, higher than few-shot Llama-3.1-70B and Qwen-2.5-72B. Encoder models outperform GPT-2, and scaling improves parity (e.g., DeBERTa-v2-XL at $\approx 0.28$ vs. DeBERTa-v3-large at 0.39 for FNR regarding {GEN,SO}). As with per-axis results, reweighting reduces FNR gaps but can slightly raise FPR gaps.
Takeaway. Detecting multi-demographic biases remains particularly difficult for LLM-based methods. Fine-tuned models achieve relatively low disparities regarding single axes but struggle with biases targeting multiple demographics. These results highlight intersectional disparities as an important direction for future research.
6.  Discussion
Our study provides a practical assessment of how LLMs can scale the detection of demographic-targeted social biases in text, supporting compliance with governance and regulatory standards. In this section, we discuss broader implications, key limitations, and directions for future work.
Inclusive stakeholder input. Automation enables large-scale auditing while minimizing human exposure to harmful content Steiger et al. (2021). However, bias detection remains a socio-technical challenge: social biases are culturally situated, and the detection pipeline must be guided by participatory processes Markov et al. (2023).
Data coverage. Our evaluation relies mainly on English datasets reflecting global North contexts, which limits generalizability to other languages and regions. Progress will require multilingual and cross-lingual benchmarks Corazza et al. (2020); Huang and Xiong (2023); Neplenbroek et al. (2024), methods adapted for low-resource settings, and attention to language-specific biases that may not align with English-centric taxonomies. Dataset diversity should also ensure sufficient coverage of social biases that may exist in different cultural contexts. Coverage should also extend to more categories, e.g., political beliefs and biases encoded across multiple data modalities.
Methodological opportunities. Our findings highlight both promises and existing limitations of scalable automated bias detection. Future work should further explore advanced prompting (e.g., chain-of-thought Wei et al. (2022)) and efficient fine-tuning (e.g., using adapters). To handle multimodal data, exploring effective and scalable solutions remains an interesting open challenge. Moreover, while our results offer initial guidance, balancing the potential gains from training and deploying larger models and the scalability required for auditing massive corpora warrants deeper exploration. Finally, our study shows that detecting infrequent, subtle, and multi-axis biases remains a crucial underexplored direction for future research.
7.  Bibliographical References
[…]
  • Barikeri et al. (2021) Soumya Barikeri, Anne Lauscher, Ivan Vulić, and Goran Glavaš. 2021. Redditbias: A real-world resource for bias evaluation and debiasing of conversational language models. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 1941–1955.
  • Bartl et al. (2020) Marion Bartl, Malvina Nissim, and Albert Gatt. 2020. Unmasking contextual stereotypes: Measuring and mitigating bert’s gender bias. In Proceedings of the Second Workshop on Gender Bias in Natural Language Processing.
  • Blodgett et al. (2020) Su Lin Blodgett, Solon Barocas, Hal Daumé III, and Hanna Wallach. 2020. Language (technology) is power: A critical survey of “bias” in NLP. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 5454–5476, Online. Association for Computational Linguistics.
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems, 33:1877–1901.
  • Chen et al. (2024a) Jianfa Chen, Emily Shen, Trupti Bavalatti, Xiaowen Lin, Yongkai Wang, Shuming Hu, Harihar Subramanyam, Ksheeraj Sai Vepuri, Ming Jiang, Ji Qi, et al. 2024a. Class-rag: Real-time content moderation with retrieval augmented generation. arXiv preprint arXiv:2410.14881.
  • Chen et al. (2024b) Jianlv Chen, Shitao Xiao, Peitian Zhang, Kun Luo, Defu Lian, and Zheng Liu. 2024b. Bge m3-embedding: Multi-lingual, multi-functionality, multi-granularity text embeddings through self-knowledge distillation. arXiv preprint arXiv:2402.03216.
  • Congress (1964) United States Congress. 1964. Title vii of the civil rights act of 1964. Pub. L. No. 88-352, 78 Stat. 241; codified at 42 U.S.C. § 2000e et seq.
  • Corazza et al. (2020) Michele Corazza, Stefano Menini, Elena Cabrio, Sara Tonelli, and Serena Villata. 2020. A multilingual evaluation for online hate speech detection. ACM Transactions on Internet Technology (TOIT), 20(2):1–22.
  • Dodge et al. (2021) Jesse Dodge, Maarten Sap, Ana Marasović, William Agnew, Gabriel Ilharco, Dirk Groeneveld, Margaret Mitchell, and Matt Gardner. 2021. Documenting large webtext corpora: A case study on the colossal clean crawled corpus. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 1286–1305, Online and Punta Cana, Dominican Republic. Association for Computational Linguistics.
  • Dubey et al. (2024) Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Amy Yang, Angela Fan, et al. 2024. The llama 3 herd of models. arXiv preprint arXiv:2407.21783.
  • ElSherief et al. (2021) Mai ElSherief, Caleb Ziems, David Muchlinski, Vaishnavi Anupindi, Jordyn Seybolt, Munmun De Choudhury, and Diyi Yang. 2021. Latent hatred: A benchmark for understanding implicit hate speech. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 345–363.
  • EU FRA (2018) EU FRA. 2018. Handbook on European Non-Discrimination Law. Publications Office of the European Union, Luxembourg.
  • European Commission (2025) European Commission. 2025. Third draft of the general-purpose ai code of practice. Draft prepared by independent experts under the coordination of the European AI Office.
  • European Union (2024) European Union. 2024. Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence. [Accessed: 2025-03-27].
  • Felkner et al. (2023) Virginia Felkner, Ho-Chun Herbert Chang, Eugene Jang, and Jonathan May. 2023. Winoqueer: A community-in-the-loop benchmark for anti-lgbtq+ bias in large language models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 9126–9140.
  • Gallegos et al. (2024) Isabel O. Gallegos, Ryan A. Rossi, Joe Barrow, Md Mehrab Tanjim, Sungchul Kim, Franck Dernoncourt, Tong Yu, Ruiyi Zhang, and Nesreen K. Ahmed. 2024. Bias and fairness in large language models: A survey. Computational Linguistics, 50(3):1097–1179.
  • Gehman et al. (2020) Samuel Gehman, Suchin Gururangan, Maarten Sap, Yejin Choi, and Noah A Smith. 2020. Realtoxicityprompts: Evaluating neural toxic degeneration in language models. arXiv preprint arXiv:2009.11462.
  • GLM et al. (2024) Team GLM, Aohan Zeng, Bin Xu, Bowen Wang, Chenhui Zhang, Da Yin, Dan Zhang, Diego Rojas, Guanyu Feng, Hanlin Zhao, et al. 2024. Chatglm: A family of large language models from glm-130b to glm-4 all tools. arXiv preprint arXiv:2406.12793.
  • Hardt et al. (2016) Moritz Hardt, Eric Price, and Nati Srebro. 2016. Equality of opportunity in supervised learning. Advances in neural information processing systems, 29.
  • Hartvigsen et al. (2022) Thomas Hartvigsen, Saadia Gabriel, Hamid Palangi, Maarten Sap, Dipankar Ray, and Ece Kamar. 2022. Toxigen: A large-scale machine-generated dataset for adversarial and implicit hate speech detection. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 3309–3326.
  • He et al. (2020) Pengcheng He, Xiaodong Liu, Jianfeng Gao, and Weizhu Chen. 2020. Deberta: Decoding-enhanced bert with disentangled attention. arXiv preprint arXiv:2006.03654.
  • Huang and Xiong (2023) Yufei Huang and Deyi Xiong. 2023. Cbbq: A chinese bias benchmark dataset curated with human-ai collaboration for large language models. arXiv preprint arXiv:2306.16244.
  • Hube and Fetahu (2018) Christoph Hube and Besnik Fetahu. 2018. Detecting biased statements in wikipedia. In Companion proceedings of the the web conference 2018, pages 1779–1786.
  • Inan et al. (2023) Hakan Inan, Kartikeya Upasani, Jianfeng Chi, Rashi Rungta, Krithika Iyer, Yuning Mao, Michael Tontchev, Qing Hu, Brian Fuller, Davide Testuggine, et al. 2023. Llama guard: Llm-based input-output safeguard for human-ai conversations. arXiv preprint arXiv:2312.06674.
  • Kamiran and Calders (2012) Faisal Kamiran and Toon Calders. 2012. Data preprocessing techniques for classification without discrimination. Knowledge and information systems, 33(1):1–33.
  • Kearns et al. (2018) Michael Kearns, Seth Neel, Aaron Roth, and Zhiwei Steven Wu. 2018. Preventing fairness gerrymandering: Auditing and learning for subgroup fairness. In International conference on machine learning, pages 2564–2572. PMLR.
  • Klose et al. (2025) Alexander Klose, Doris Liebscher, Maria Wersig, and Michael Wrase, editors. 2025. Landesantidiskriminierungsgesetz Berlin. Nomos, Germany.
  • Koo et al. (2024) Ryan Koo, Minhwa Lee, Vipul Raheja, Jong Inn Park, Zae Myung Kim, and Dongyeop Kang. 2024. Benchmarking cognitive biases in large language models as evaluators. In Findings of the Association for Computational Linguistics ACL 2024, pages 517–545.
  • Kreutzer et al. (2022) Julia Kreutzer, Isaac Caswell, Lisa Wang, Ahsan Wahab, Daan Van Esch, Nasanbayar Ulzii-Orshikh, Allahsera Tapo, Nishant Subramani, Artem Sokolov, Claytone Sikasote, et al. 2022. Quality at a glance: An audit of web-crawled multilingual datasets. Transactions of the Association for Computational Linguistics, 10:50–72.
  • Kumar et al. (2024) Deepak Kumar, Yousef Anees AbuHashem, and Zakir Durumeric. 2024. Watch your language: Investigating content moderation with large language models. In Proceedings of the International AAAI Conference on Web and Social Media, volume 18, pages 865–878.
  • Li et al. (2023) Yunqi Li, Lanjing Zhang, and Yongfeng Zhang. 2023. Fairness of chatgpt. arXiv preprint arXiv:2305.18569.
  • Litt (1961) Edgar Litt. 1961. Jewish ethno-religious involvement and political liberalism. Social Forces, 39(4):328–332.
  • Liu (2019) Yinhan Liu. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692, 364.
  • Luccioni and Viviano (2021) Alexandra Luccioni and Joseph Viviano. 2021. What’s in the box? an analysis of undesirable content in the Common Crawl corpus. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 2: Short Papers), pages 182–189, Online. Association for Computational Linguistics.
  • Maheshwari et al. (2024) Gaurav Maheshwari, Dmitry Ivanov, and Kevin El Haddad. 2024. Efficacy of synthetic data as a benchmark. arXiv preprint arXiv:2409.11968.
  • Markov et al. (2023) Todor Markov, Chong Zhang, Sandhini Agarwal, Florentine Eloundou Nekoul, Theodore Lee, Steven Adler, Angela Jiang, and Lilian Weng. 2023. A holistic approach to undesired content detection in the real world. In Proceedings of the AAAI Conference on Artificial Intelligence, volume 37, pages 15009–15018.
  • Mathew et al. (2021) Binny Mathew, Punyajoy Saha, Seid Muhie Yimam, Chris Biemann, Pawan Goyal, and Animesh Mukherjee. 2021. Hatexplain: A benchmark dataset for explainable hate speech detection. In Proceedings of the AAAI conference on artificial intelligence, volume 35, pages 14867–14875.
  • Nadeem et al. (2021) Moin Nadeem, Anna Bethke, and Siva Reddy. 2021. Stereoset: Measuring stereotypical bias in pretrained language models. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 5356–5371.
  • Nangia et al. (2020) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel Bowman. 2020. Crows-pairs: A challenge dataset for measuring social biases in masked language models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 1953–1967.
  • Navigli et al. (2023) Roberto Navigli, Simone Conia, and Björn Ross. 2023. Biases in large language models: origins, inventory, and discussion. ACM Journal of Data and Information Quality, 15(2):1–21.
  • Neplenbroek et al. (2024) Vera Neplenbroek, Arianna Bisazza, and Raquel Fernández. 2024. Mbbq: A dataset for cross-lingual comparison of stereotypes in generative llms. arXiv preprint arXiv:2406.07243.
  • Palla et al. (2025) Konstantina Palla, José Luis Redondo García, Claudia Hauff, Francesco Fabbri, Andreas Damianou, Henrik Lindström, Dan Taber, and Mounia Lalmas. 2025. Policy-as-prompt: Rethinking content moderation in the age of large language models. In Proceedings of the 2025 ACM Conference on Fairness, Accountability, and Transparency, pages 840–854.
  • Parrish et al. (2022) Alicia Parrish, Angelica Chen, Nikita Nangia, Vishakh Padmakumar, Jason Phang, Jana Thompson, Phu Mon Htut, and Samuel Bowman. 2022. Bbq: A hand-built bias benchmark for question answering. In Findings of the Association for Computational Linguistics: ACL 2022.
  • Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog, 1(8):9.
  • Raza et al. (2024) Shaina Raza, Muskan Garg, Deepak John Reji, Syed Raza Bashir, and Chen Ding. 2024. Nbias: A natural language processing framework for bias identification in text. Expert Systems with Applications, 237:121542.
  • Rudinger et al. (2018) Rachel Rudinger, Jason Naradowsky, Brian Leonard, and Benjamin Van Durme. 2018. Gender bias in coreference resolution. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), pages 8–14.
  • Sap et al. (2020) Maarten Sap, Saadia Gabriel, Lianhui Qin, Dan Jurafsky, Noah A Smith, and Yejin Choi. 2020. Social bias frames: Reasoning about social and power implications of language. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 5477–5490.
  • Sorower (2010) Mohammad S Sorower. 2010. A literature survey on algorithms for multi-label learning. Oregon State University, Corvallis, 18(1):25.
  • Steiger et al. (2021) Miriah Steiger, Timir J Bharucha, Sukrit Venkatagiri, Martin J Riedl, and Matthew Lease. 2021. The psychological well-being of content moderators: the emotional labor of commercial moderation and avenues for improving support. In Proceedings of the 2021 CHI conference on human factors in computing systems, pages 1–14.
  • Sun et al. (2024) Lichao Sun, Yue Huang, Haoran Wang, Siyuan Wu, Qihui Zhang, Chujie Gao, Yixin Huang, Wenhan Lyu, Yixuan Zhang, Xiner Li, et al. 2024. Trustllm: Trustworthiness in large language models. arXiv preprint arXiv:2401.05561, 3.
  • Tabassi (2023) Elham Tabassi. 2023. Artificial intelligence risk management framework (ai rmf 1.0).
  • Vashney (2022) Kush R Vashney. 2022. Trustworthy machine learning. Independently published.
  • Viprey (2002) Mouna Viprey. 2002. New anti-discrimination law adopted. Eurofound. Published: 3 January 2002.
  • Wang et al. (2023) Boxin Wang, Weixin Chen, Hengzhi Pei, Chulin Xie, Mintong Kang, Chenhui Zhang, Chejian Xu, Zidi Xiong, Ritik Dutta, Rylan Schaeffer, et al. 2023. Decodingtrust: A comprehensive assessment of trustworthiness in gpt models. In NeurIPS.
  • Wang et al. (2024) Song Wang, Peng Wang, Tong Zhou, Yushun Dong, Zhen Tan, and Jundong Li. 2024. Ceb: Compositional evaluation benchmark for fairness in large language models. arXiv preprint arXiv:2407.02408.
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems, 35:24824–24837.
  • Wolf et al. (2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander M. Rush. 2020. Transformers: State-of-the-art natural language processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 38–45, Online. Association for Computational Linguistics.
  • Yang et al. (2024) An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, et al. 2024. Qwen2. 5 technical report. arXiv preprint arXiv:2412.15115.
  • Youdao (2023) NetEase Youdao. 2023. Bcembedding: Bilingual and crosslingual embedding for rag. https://github.com/netease-youdao/BCEmbedding.
  • Zafar et al. (2017) Muhammad Bilal Zafar, Isabel Valera, Manuel Gomez Rodriguez, and Krishna P Gummadi. 2017. Fairness beyond disparate treatment & disparate impact: Learning classification without disparate mistreatment. In Proceedings of the 26th international conference on world wide web, pages 1171–1180.
  • Zeng et al. (2024) Wenjun Zeng, Yuchi Liu, Ryan Mullins, Ludovic Peran, Joe Fernandez, Hamza Harkous, Karthik Narasimhan, Drew Proud, Piyush Kumar, Bhaktipriya Radharapu, et al. 2024. Shieldgemma: Generative ai content moderation based on gemma. arXiv preprint arXiv:2407.21772.
  • Zhan et al. (2025) Xianyang Zhan, Agam Goyal, Yilun Chen, Eshwar Chandrasekharan, and Koustuv Saha. 2025. Slm-mod: Small language models surpass llms at content moderation. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 8774–8790.
  • Zhao et al. (2018) Jieyu Zhao, Tianlu Wang, Mark Yatskar, Vicente Ordonez, and Kai-Wei Chang. 2018. Gender bias in coreference resolution: Evaluation and debiasing methods. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), pages 15–20.
Appendix A Data characteristics
Figure 1: Token count distribution
Figure 2: Social bias label co-occurrence.
(a) Binary biased-unbiased counts (train subset)
(b) Number of labels per text (train subset)
Figure 3: Social bias label distribution
Text statistics. In Figure 1, we show the token count distribution for all the bias categories, including the unbiased texts. To obtain tokens, we use the RoBERTa tokenizer. We see that the token counts for each text instance are generally low, indicating our dataset mainly contains short text instances. However, there are some outliers, e.g., in GEN and SO, where we see around 400 token length texts. Upon inspection, we observed that such instances primarily come from the BBQ dataset, that have very long context texts.
Label statistics. In Figure 3, we visualize the label statistics in the final curated dataset. The visualizations show label imbalances in the data, highlighting the need for weighted loss for optimization and motivating future work to explore further fairness interventions to ensure equitable bias detection performance. The statistics show that our data contains more biased instances than unbiased ones. Furthermore, we see that most instances target a single demographic axis. However, many instances target two axes. Instances targeting more than two demographic axes are significantly fewer in our dataset. We provide more detailed label co-occurrence statistics in Figure 2. The figure shows that text instances target specific demographics more often. For instance, texts target RAC and GEN more often. Similarly, texts target DIS,AGE, and PHY relatively less often. Furthermore, we see that GEN co-occurs with many other demographic axes, e.g., SO, RAC, and DIS. Note that while RAC and REL appear together frequently, many of these instances simply target “Jewish identities.”
Figure 4: Prompt used for LLMs to detect demographic-targeted biases
Figure 5: Prompting pipeline adapted for social bias detection
Appendix B Practical setup of testbed
B.1.  Prompting
All LLMs are accessed via API through an MLOps platform. We fix temperature to 0 and top_p to 1, ensuring deterministic outputs by selecting the model’s most likely generation while still allowing consideration of the full token space. For in-context learning, we embed the training and development sets using BGE-M3 or BCEmbedding models. At inference time, we compute cosine similarity between the query and development set vectors to retrieve the top-$K$ few-shot examples. As a baseline, we also apply a random few-shot selection from the training set, with balanced sampling between biased and unbiased texts. Model predictions are extracted via pattern matching. Responses that cannot be parsed through pattern matching to assign to one of the multiple demographic axes are marked as “invalid.” The social bias policy used in the text prompt is shown here.
S1: Gender and Sexual Identity Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content based on gender and sexual identity. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - stereotypes on gender roles and gender professions ("The woman is cooking") - toxic and hateful text targeting gender group or sexual identity SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S2: Sexual Orientation Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content based on sexual orientation. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting sexual orientation SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S3: Disability Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content targeting physically or mentally disabled people. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting people with disability SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S4: Age Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content targeting people based on age like children, young people, and old people. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting people of different age groups SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S5: Race and Ethnicity Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content targeting people based on race, ethnicity, skin color, or minorities. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting races and ethnicities SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S6: Nationality Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content targeting people based on their nationality or target a specific country. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting a nation or people from the nation SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S7: Religion Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content targeting a religion or people holding certain religious beliefs. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting a religion SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S8: Socioeconomic Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content targeting people based on their socioeconomic status like targeting people in poverty, targeting rich people, targeting the homeless, etc. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting socioeconomic groups SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S9: Physical and Body Appearance Bias. The text should not perpetuate commonly held stereotypes, toxic, or hateful content targeting people based on their physical appearance or body structure like fat or obese people. BIAS EXAMPLES: - stereotypes that attribute simplistic positive or negative traits - toxic and hateful text targeting physical appearance SAFE CASES: - contradict common stereotypes or anti-stereotypes - discuss, question, and challenge stereotype, toxic, or hate speech S10: Safe and Unbiased Text. The text does not fall into any of the other categories. **DO NOT SELECT THIS CLASS IF TEXT BELONGS TO ANY OTHER CATEGORY.** REASONS FOR SAFE: - No common stereotype description used against demographics - Anti-stereotype description used for demographics ("The woman is physically strong") - Gender role or gender occupation association is opposite of common stereotype - No violence, hate, toxic content.
B.2.  Fine-tuning
We fine-tune LLMs for sequence classification using HuggingFace’s transformers library Wolf et al. (2020), with a maximum input length of 512 tokens. For GPT-2 models, sequences are left-padded with the EOS token.
Optimization uses AdamW with linear learning rate decay, weight decay of 0.01, and gradient clipping at 1.0. To address class imbalance, we experiment with reweighted binary cross-entropy loss, where weights are derived from label frequencies in the training set. Models are trained for four epochs without reweighting and six epochs with reweighting. The effective batch size is fixed at 32, with gradient accumulation applied for larger models. Learning rates are tuned by monitoring validation loss. For each model, we use the following learning rates for optimization:
  • $10^{-4}$ (GPT-2-XL)
  • $5\times 10^{-5}$ (GPT-2-large)
  • $2\times 10^{-5}$ (RoBERTa-base)
  • $10^{-5}$ (RoBERTa-large, DeBERTa-v2-XL)
  • $5\times 10^{-6}$ (DeBERTa-v3-large)
Learning rates are not changed across loss functions (default or reweighted). Training is performed in float32 precision, except GPT-2-XL, which uses bfloat16. All experiments run on a single GPU with 32GB VRAM and 128GB host memory.
Appendix C Additional evaluations
Table 6: Analyzing the importance of retrieval augmented (RAG-based using BGE-M3) example selection for few-shot prompting by comparing performance to random sampling.
Model Setup Few-shot Binary Prediction Multi-label Prediction
$F_{1}$ FPR FNR MR HL $F_{1}^{\mu}$ $F_{1}^{\text{M}}$
Llama-Guard-8B Random 5 $66.97_{\pm 0.76}$ $0.152_{\pm 0.011}$ $0.470_{\pm 0.008}$ $339_{\pm 0.008} 0.0$$089_{\pm 0.001}$ $51.39_{\pm 0.80}$ $33.39_{\pm 1.62}$
10 $65.55_{\pm 0.77}$ $0.147_{\pm 0.011}$ $0.488_{\pm 0.009}$ $288_{\pm 0.008} 0.0$$099_{\pm 0.001}$ $45.65_{\pm 0.85}$ $30.81_{\pm 1.67}$
RAG 5 $75.16_{\pm 0.64}$ $0.192_{\pm 0.012}$ $0.358_{\pm 0.009}$ $485_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.66_{\pm 0.68}$ $46.24_{\pm 1.87}$
10 $75.17_{\pm 0.64}$ $0.186_{\pm 0.011}$ $0.359_{\pm 0.008}$ $486_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.79_{\pm 0.69}$ $44.68_{\pm 1.82}$
Llama-3.1-8B Random 5 $84.50_{\pm 0.45}$ $832_{\pm 0.011} 0.0$$57_{\pm 0.004} 0.0$$075_{\pm 0.004}$ $0.236_{\pm 0.004}$ $47.73_{\pm 0.49}$ $33.74_{\pm 0.49}$
10 $84.19_{\pm 0.46}$ $698_{\pm 0.014} 0.0$$97_{\pm 0.005} 0.0$$051_{\pm 0.004}$ $0.252_{\pm 0.004}$ $45.21_{\pm 0.48}$ $33.38_{\pm 0.43}$
RAG 5 $87.27_{\pm 0.40}$ $752_{\pm 0.012} 0.0$$023_{\pm 0.003}$ $0.411_{\pm 0.008}$ $0.140_{\pm 0.004}$ $62.19_{\pm 0.68}$ $44.58_{\pm 0.73}$
10 $87.47_{\pm 0.40}$ $746_{\pm 0.013} 0.0$$021_{\pm 0.002}$ $0.501_{\pm 0.009}$ $0.127_{\pm 0.004}$ $64.69_{\pm 0.70}$ $45.96_{\pm 0.82}$
GLM-4-9B Random 5 $83.81_{\pm 0.46}$ $783_{\pm 0.011} 0.0$$082_{\pm 0.005}$ $457_{\pm 0.009} 0.0$$095_{\pm 0.002}$ $63.37_{\pm 0.71}$ $51.23_{\pm 1.57}$
10 $83.65_{\pm 0.47}$ $761_{\pm 0.012} 0.0$$091_{\pm 0.005}$ $475_{\pm 0.008} 0.0$$090_{\pm 0.002}$ $64.69_{\pm 0.67}$ $52.79_{\pm 1.56}$
RAG 5 $87.10_{\pm 0.40}$ $774_{\pm 0.012} 0.0$$021_{\pm 0.003}$ $773_{\pm 0.007} 0.0$$036_{\pm 0.001}$ $85.95_{\pm 0.50}$ $73.43_{\pm 1.69}$
10 $86.98_{\pm 0.41}$ $775_{\pm 0.012} 0.0$$023_{\pm 0.003}$ $782_{\pm 0.007} 0.0$$034_{\pm 0.001}$ $86.74_{\pm 0.48}$ $75.46_{\pm 1.68}$
Llama-3.1-70B Random 5 $84.28_{\pm 0.47}$ $0.541_{\pm 0.014}$ $0.134_{\pm 0.006}$ $284_{\pm 0.008} 0.0$$095_{\pm 0.001}$ $67.96_{\pm 0.45}$ $58.86_{\pm 1.54}$
10 $84.01_{\pm 0.46}$ $0.511_{\pm 0.014}$ $0.147_{\pm 0.006}$ $289_{\pm 0.008} 0.0$$098_{\pm 0.001}$ $66.29_{\pm 0.47}$ $56.95_{\pm 1.49}$
RAG 5 $88.49_{\pm 0.38}$ $581_{\pm 0.014} 0.0$$046_{\pm 0.004}$ $657_{\pm 0.008} 0.0$$046_{\pm 0.001}$ $83.28_{\pm 0.42}$ $73.16_{\pm 1.36}$
10 $88.82_{\pm 0.39}$ $557_{\pm 0.015} 0.0$$046_{\pm 0.004}$ $648_{\pm 0.008} 0.0$$047_{\pm 0.001}$ $83.08_{\pm 0.41}$ $75.07_{\pm 1.39}$
Qwen-2.5-72B Random 5 $82.02_{\pm 0.51}$ $0.638_{\pm 0.014}$ $0.151_{\pm 0.006}$ $0.208_{\pm 0.007}$ $0.135_{\pm 0.002}$ $58.98_{\pm 0.54}$ $44.85_{\pm 0.79}$
10 $80.81_{\pm 0.51}$ $0.600_{\pm 0.014}$ $0.181_{\pm 0.007}$ $0.177_{\pm 0.007}$ $0.143_{\pm 0.002}$ $55.87_{\pm 0.54}$ $43.85_{\pm 0.80}$
RAG 5 $87.24_{\pm 0.39}$ $551_{\pm 0.014} 0.0$$078_{\pm 0.005}$ $583_{\pm 0.008} 0.0$$065_{\pm 0.002}$ $77.33_{\pm 0.55}$ $60.44_{\pm 1.15}$
10 $87.38_{\pm 0.41}$ $552_{\pm 0.014} 0.0$$075_{\pm 0.004}$ $600_{\pm 0.009} 0.0$$060_{\pm 0.002}$ $78.94_{\pm 0.52}$ $63.00_{\pm 1.19}$
C.1.  Ablation study: in-context learning
We evaluate the impact of retrieval-augmented generation (RAG) on few-shot example selection compared to random sampling. Results are presented in Table 6. Overall, RAG consistently enhances bias detection performance.
In binary classification, RAG achieves higher $F_{1}$ scores across all models. Improvements in detection metrics are consistent across model sizes, demonstrating the benefit of providing LLMs with semantically similar examples during in-context learning. RAG generally reduces False Negative Rates (FNR), though it occasionally causes slight increases in False Positive Rates (FPR), as observed with Llama Guard-3-8B and GLM-4-9B. This tradeoff is typically favorable, since reducing FNR is crucial for minimizing missed detections. Notably, while adding more examples under RAG yields only modest additional gains, increasing the number of randomly selected examples often leads to degraded performance.
For multi-label prediction, RAG delivers even greater improvements over random sampling. As in the binary case, providing more RAG-selected examples enhances performance, whereas adding more random examples consistently worsens detection outcomes. This highlights an important insight: supplying more relevant examples benefits prompting-based detection, but including irrelevant examples can be detrimental.
In summary, RAG significantly strengthens in-context learning by providing more meaningful examples, resulting in higher accuracy and improved multi-label predictions. Although small increases in FPR can occur, the overall gains clearly favor RAG over random sampling.
Table 7: Comparing few-shot prompting performance when using different retrieval embedding models BGE-M3 and BCEmbedding (BCEmb.).
Model Setup Few-shot Binary Prediction Multi-label Prediction
$F_{1}$ FPR FNR MR HL $F_{1}^{\mu}$ $F_{1}^{\text{M}}$
Llama-Guard-8B BGE-M3 5 $75.16_{\pm 0.64}$ $0.192_{\pm 0.012}$ $0.358_{\pm 0.009}$ $485_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.66_{\pm 0.68}$ $46.24_{\pm 1.87}$
10 $75.17_{\pm 0.64}$ $0.186_{\pm 0.011}$ $0.359_{\pm 0.008}$ $486_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.79_{\pm 0.69}$ $44.68_{\pm 1.82}$
BCEmb. 5 $73.90_{\pm 0.66}$ $0.196_{\pm 0.011}$ $0.374_{\pm 0.008}$ $478_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.32_{\pm 0.68}$ $44.21_{\pm 1.78}$
10 $74.07_{\pm 0.66}$ $0.184_{\pm 0.012}$ $0.374_{\pm 0.009}$ $482_{\pm 0.008} 0.0$$067_{\pm 0.001}$ $65.60_{\pm 0.70}$ $42.99_{\pm 1.83}$
Llama-3.1-8B BGE-M3 5 $87.27_{\pm 0.40}$ $752_{\pm 0.012} 0.0$$023_{\pm 0.003}$ $0.411_{\pm 0.008}$ $0.140_{\pm 0.004}$ $62.19_{\pm 0.68}$ $44.58_{\pm 0.73}$
10 $87.47_{\pm 0.40}$ $746_{\pm 0.013} 0.0$$021_{\pm 0.002}$ $0.501_{\pm 0.009}$ $0.127_{\pm 0.004}$ $64.69_{\pm 0.70}$ $45.96_{\pm 0.82}$
BCEmb. 5 $87.18_{\pm 0.38}$ $750_{\pm 0.013} 0.0$$026_{\pm 0.003}$ $0.464_{\pm 0.008}$ $0.125_{\pm 0.004}$ $64.84_{\pm 0.68}$ $46.36_{\pm 0.76}$
10 $87.46_{\pm 0.41}$ $740_{\pm 0.013} 0.0$$023_{\pm 0.003}$ $0.552_{\pm 0.008}$ $0.113_{\pm 0.004}$ $67.22_{\pm 0.72}$ $47.69_{\pm 0.85}$
GLM-4-9B BGE-M3 5 $87.10_{\pm 0.40}$ $774_{\pm 0.012} 0.0$$021_{\pm 0.003}$ $773_{\pm 0.007} 0.0$$036_{\pm 0.001}$ $85.95_{\pm 0.50}$ $73.43_{\pm 1.69}$
10 $86.98_{\pm 0.41}$ $775_{\pm 0.012} 0.0$$023_{\pm 0.003}$ $782_{\pm 0.007} 0.0$$034_{\pm 0.001}$ $86.74_{\pm 0.48}$ $75.46_{\pm 1.68}$
BCEmb. 5 $86.79_{\pm 0.41}$ $783_{\pm 0.012} 0.0$$025_{\pm 0.003}$ $802_{\pm 0.007} 0.0$$032_{\pm 0.001}$ $87.50_{\pm 0.48}$ $74.80_{\pm 1.71}$
10 $86.95_{\pm 0.43}$ $769_{\pm 0.012} 0.0$$025_{\pm 0.003}$ $808_{\pm 0.007} 0.0$$031_{\pm 0.001}$ $87.90_{\pm 0.47}$ $75.27_{\pm 1.66}$
Llama-3.1-70B BGE-M3 5 $88.49_{\pm 0.38}$ $581_{\pm 0.014} 0.0$$046_{\pm 0.004}$ $657_{\pm 0.008} 0.0$$046_{\pm 0.001}$ $83.28_{\pm 0.42}$ $73.16_{\pm 1.36}$
10 $88.82_{\pm 0.39}$ $557_{\pm 0.015} 0.0$$046_{\pm 0.004}$ $648_{\pm 0.008} 0.0$$047_{\pm 0.001}$ $83.08_{\pm 0.41}$ $75.07_{\pm 1.39}$
BCEmb. 5 $88.41_{\pm 0.39}$ $577_{\pm 0.015} 0.0$$049_{\pm 0.004}$ $692_{\pm 0.008} 0.0$$041_{\pm 0.001}$ $84.63_{\pm 0.41}$ $74.11_{\pm 1.51}$
10 $88.75_{\pm 0.39}$ $546_{\pm 0.015} 0.0$$051_{\pm 0.004}$ $693_{\pm 0.008} 0.0$$041_{\pm 0.001}$ $84.80_{\pm 0.41}$ $76.60_{\pm 1.61}$
Qwen-2.5-72B BGE-M3 5 $87.24_{\pm 0.39}$ $551_{\pm 0.014} 0.0$$078_{\pm 0.005}$ $583_{\pm 0.008} 0.0$$065_{\pm 0.002}$ $77.33_{\pm 0.55}$ $60.44_{\pm 1.15}$
10 $87.38_{\pm 0.41}$ $552_{\pm 0.014} 0.0$$075_{\pm 0.004}$ $600_{\pm 0.009} 0.0$$060_{\pm 0.002}$ $78.94_{\pm 0.52}$ $63.00_{\pm 1.19}$
BCEmb. 5 $86.76_{\pm 0.44}$ $565_{\pm 0.014} 0.0$$083_{\pm 0.005}$ $617_{\pm 0.009} 0.0$$060_{\pm 0.002}$ $78.54_{\pm 0.56}$ $60.96_{\pm 1.22}$
10 $87.25_{\pm 0.41}$ $557_{\pm 0.014} 0.0$$076_{\pm 0.004}$ $638_{\pm 0.008} 0.0$$054_{\pm 0.002}$ $80.42_{\pm 0.52}$ $64.07_{\pm 1.29}$
C.2.  Ablation study: Embedding model
We next examine how the choice of embedding model affects in-context learning performance for prompting, comparing BGE-M3 Chen et al. (2024b) and BCEmbedding Youdao (2023) for selecting in-context examples. The results are presented in Table 7.
BGE-M3 exhibits a slight but consistent advantage in binary bias detection, producing marginally higher $F_{1}$ scores across multiple LLMs. However, the overall differences are minimal. In contrast, for multi-label prediction, BCEmbedding performs slightly better on metrics such as MR for many models. This finding suggests that while both embedding models select examples that yield similar overall outcomes, subtle differences exist. Specifically, BGE-M3-selected examples tend to improve binary bias detection by helping models better distinguish biased from unbiased samples, whereas BCEmbedding-selected examples slightly enhance the detection of specific bias types within biased instances.
Overall, both embedding models deliver strong and comparable performance for in-context learning, with only minor trade-offs. Their results indicate that either embedding model is well-suited for bias detection tasks.
Figure 6: Per-label $F_{1}$ scores for prompting. Models used 10-shot RAG-based in-context using BGE-M3.
Figure 7: Per-label $F_{1}$ scores for fine-tuning different models using default loss or reweighted (bal) loss.
C.3.  Bias detection of each bias class
We now analyze model performance across different demographic targets. Specifically, we examine the $F_{1}$ scores for all demographic axes in our taxonomy that may be subject to bias. Figure 6 presents results for various prompted models using 10-shot RAG-based in-context learning with BGE-M3, while Figure 7 compares fine-tuned models and explores the impact of reweighted loss (“bal” in the figure) across demographics.
Our analysis shows that fine-tuned models consistently outperform prompting and transfer learning across all bias classes. The most notable $F_{1}$ score gains appear in the AGE and SES categories, which are less frequent in the dataset.
Among the prompted LLMs, Llama-3.1-70B achieves the highest $F_{1}$ scores across nearly all bias categories, except for GEN, where GLM-9B—despite being much smaller—slightly outperforms it. Interestingly, Qwen-2.5-72B, though the largest LLM, performs worse in many low-frequency categories such as DIS, AGE, and SES. It performs comparably to the best prompting models only for GEN and RAC, which are the most common categories in the benchmark.
For fine-tuned models, encoder-only architectures (e.g., RoBERTa and DeBERTa) generally outperform decoder-only language models, i.e., GPT-2, across most demographic axes. The trends mirror those observed in the prompting setup: models achieve their best performance for SES, while NAT consistently shows the lowest $F_{1}$ scores. Reweighted loss often improves detection performance or yields similar results to the default loss. For example, in NAT, the axis that suffered from the weakest detection performance, reweighted loss improves performance across all models. However, improvements are not universal. For instance, GPT-2-large experiences slight declines in $F_{1}$ for some demographics such as AGE and SES when reweighted loss is applied.
These findings provide additional insight into the disparity results discussed in Section 5, which highlight performance gaps across demographic axes. This deeper analysis underscores the need to develop more nuanced methods that can mitigate detection disparities without substantially compromising overall performance.